OpenStudy (anonymous):

I don't get it if i have to mucltiply the exoponents or just add them use the product rule to simplify the expression. write answer in exponential form. (-6)^4*(-6)^4

OpenStudy (amistre64):

lets start at what an exponent means: a^3 means a*a*a so lets spose we have a^3 * a^3 that tells us simply: (a*a*a) * (a*a*a) = a*a*a*a*a*a and we can rewrite it back into exponential notation as a^6 from this we can conclude that: a^3 * a^3 = a^(3+3) = a^6

OpenStudy (anonymous):

so it is 6^8 I just add the exponents

OpenStudy (amistre64):

yep; but they might want it as (-6)^8

OpenStudy (anonymous):

thanks a lot:)

OpenStudy (amistre64):

youre welcome